# LLMs can Perform Multi-Dimensional Analytic Writing Assessments: A Case Study of L2 Graduate-Level Academic English Writing

### v1

@@ Abstract
The paper explores the performance of LLMs in the context of multi-dimensional analytic writing assessments, i.e. their ability to provide both scores and comments based on multiple assessment criteria. Using a corpus of literature reviews written by L2 graduate students and assessed by human experts against 9 analytic criteria, we prompt several popular LLMs to perform the same task under various conditions. To evaluate the quality of feedback comments, we apply a novel feedback comment quality evaluation framework. This framework is interpretable, cost-efficient, scalable, and reproducible, compared to existing methods that rely on manual judgments. We find that LLMs can generate reasonably good and generally reliable multi-dimensional analytic assessments. We release our corpus [^1] https://github.com/jaaack-wang/multi-dimensional-analytic-writing-assessments. for reproducibility.
@@ A Case Study of L2 Graduate-Level Academic English Writing
Zhengxiang Wang1thanks: Zhengxiang Wang was a research assistant at the University of Saskatchewan for the research project that led to the creation of the corpus released in this study.   Veronika Makarova2   Zhi Li2   Jordan Kodner1   Owen Rambow1 1Department of Linguistics & IACS, Stony Brook University, USA 2Department of Linguistics, University of Saskatchewan, Canada zhengxiang.wang@stonybrook.edu
@@ 1 Introduction
Figure 1: Multi-dimensional analytic assessments, where each assessment contains a score and a comment.
In recent years, large language models (LLMs) have emerged as promising tools for self-regulated writing assessments among L2 learners. A growing number of studies (Chiang and Lee, 2023; Mizumoto and Eguchi, 2023; Han et al., 2024; Yancey et al., 2023, i.a.) have indicated the general usefulness of LLMs for automated writing assessments. Given their increasing use for this task, the following question remains understudied: can LLMs provide reasonably good multi-dimensional analytic writing assessments? We use the phrase “reasonably good” intentionally, given the open-ended nature of the task, particularly generating essay-level feedback comments.
@@
    We provide comprehensive and reproducible evidence that LLMs can generate reasonably good and generally reliable multi-dimensional analytic writing assessments. This is the primary goal of this study; we do not argue in favor of a specific LLM, nor do we advocate replacing humans with LLMs for this task.
@@
    We propose and validate a novel LLM-based framework for evaluating the quality of feedback comments. This framework is time- and cost-efficient, scalable, and reproducible, compared to manual judgments. It is also interpretable, compared to direct quality ratings.
@@ 2 Related Work
Unlike prior AWE systems, LLMs can be prompted in natural language to jointly score and comment on a given essay. A growing body of research has explored the use of LLMs for assessing L2 writing. For AES, LLMs have been examined for holistic scoring (Mizumoto and Eguchi, 2023; Yancey et al., 2023; Wang and Gayed, 2024), discourse coherence scoring (Naismith et al., 2023), and multi-dimensional analytic scoring (Yavuz et al., 2024; Banno et al., 2024). For feedback comment generation, LLMs have been studied for generating corrective feedback (Mizumoto et al., 2024; Song et al., 2024), holistic feedback (Behzad et al., 2024a, b), and multi-dimensional analytic feedback (Guo and Wang, 2024; Behzad et al., 2024a; Han et al., 2024). Stahl et al. (2024) is the only study we know of which explores LLMs jointly performing scoring and feedback comment generation, but holistically. Moreover, their corpus contains short essays by native speakers and has no human reference comments.
@@ 3 Corpus
The corpus is a result of a large research project conducted at the University of Saskatchewan in 2021 with an aim to examine the developmental trajectory of literature review writing skills among L2 graduate students. The project involved three rounds of a 5-unit online tutorial series conducted over the course of 2021, with each round lasting 13 weeks (see Appendix A for details). Participation was voluntary, with 31 participants completing all five writing tasks across all rounds, and 20 further students completing at least one task before withdrawing.
@@
The corpus was used in our previous studies (Li et al., 2023a, b; Makarova et al., 2024), but has never been made public. These three studies all use only a subset of the corpus, namely essays written individually or those based on topics 1, 3, and 5.
More concretely, Li et al. (2023a, b) focus on the individual writing skill development, but without examining the feedback comments provided in the corpus. In other words, these two studies belong to the field of English for Academic Purposes, but has less relevance to AWE.
While Makarova et al. (2024) investigate whether ChatGPT can assess L2 academic English writing, they do not compare human- and ChatGPT-generated scores and comments on the basis of each assessment criterion. Rather, they simply compare ChatGPT-generated scores and comments with average scores and concatenated comments produced by multiple human assessors for all criteria, which is not only less nuanced but also over-simplistic. Moreover, their analysis of feedback comments utilizes surface-level linguistic features such as word counts, type-token ratio, comment length, and the experiments do not consider possible user-LLM interaction modes nor prompt variations, which we do in Sections 5 and  6, respectively.
In short, this study not only presents a more comprehensive and thorough evaluation of LLMs using the full corpus, instead of its subset, but also employs a different set of evaluation methodologies. As a result, we identify no substantial overlap between this study and our previous studies.
@@
The corpus authors comprise a diverse group of L2 learners, representing a wide range of first languages and enrolled in graduate programs across various disciplines at multiple Canadian universities. Their English proficiency ranged from upper-intermediate to advanced, with an average score equivalent to IELTS Band 7 based on conversions from various standardized English language tests. Scores varied from IELTS 6.5 to 8.5, with a standard deviation of 0.55.
@@
Most essays in the corpus were assessed by three (94.3%) or two (5.0%) independent human experts. As illustrated in Fig. 1, the assessments consist of scores on a 10-point scale and comments based on 9 analytic assessment criteria. While scores were required, comments were optional for the assessors. A total of six assessors with professional experience in English language teaching participated at different stages of the research project. Table 1 provides basic information about them.
The 9 assessment criteria (see Appendix A.3 for details) include: (C1) material selection; (C2) material integration and citation; (C3) quality of key components; (C4) logic of structure; (C5) content and clarity of ideas; (C6) coherence (flow of ideas) ; (C7) cohesion (use of connectors); (C8) grammar and sentence structure; and (C9) academic vocabulary.
@@ F Graduate RA 2, 3 1-5 106
4 A Novel Feedback Comment Quality Evaluation Framework
@@
To address the issues of manual judgment, we propose an automatic evaluation framework that evaluates the quality of a feedback comment in terms of its ability to effectively identify relevant writing problems within the assessed essay. As illustrated in Fig. 2 (left), the framework utilizes LLMs to extract problems identified in assessment comments and to characterize their specificity and potential helpfulness. The framework consists of the following three steps. [^3] See Appendix B for additional details and explanations.
@@
The extracted problems are further characterized along three dimensions: whether an extracted problem (1) points to a specific part of the essay, (2) includes any form of suggestion (general or specific), and (3) provides a concrete correction that can be directly applied to fix an identified problem. These classifications offer a way to assess the specificity and potential helpfulness of related comments.
@@
The results show that both human- and LLM-provided corrections are highly relevant, with answers to those three questions being “Yes” typically above 90% time (see Appendix B.3). We thus focus on the Problem Classification results when comparing human- and LLM-identified problems in the subsequent sections.
Validating the Proposed Framework
The first author and a paid graduate student in Linguistics (native speaker) first annotated some held-out samples for training and developing the annotation guidelines. Each then independently annotated at least another 200 samples containing human- and LLM-generated comments or problems for Problem Extraction and Problem Classification. Afterward, they met to resolve disagreements before the inter-annotator agreement (IAA) was calculated.
We measure IAA using Cohen’s Kappa. As is known Feinstein and Cicchetti (1990), Cohen’s Kappa can provide misleading values with highly imbalanced class distributions. We therefore also provide exact match rates which have not been corrected for random agreement. Fig. 2 (right) shows that the IAA is typically high. When the Cohen’s Kappa is low due to class imbalance (i.e., problems being incorrectly or not extracted is uncommon or rare and nearly all extracted problems contain a suggestion), the exact match rates are high. LLM task performance, evaluated based on the resolved annotations, is also notably high.
@@ 5.2 Baselines
Given the open-ended nature of the task, we compare raw assessments produced across individual assessors to understand the assessment patterns and behaviors of humans and LLMs. For a more robust statistical analysis, we only consider raw assessments made by assessors B, C, and F, since the essays they each assessed and co-assessed both cover at least half of the corpus.
@@ 5.3 Evaluation of Scores
AAR measures the percentage of scores (from two raters) that lie within a specified threshold $k$ of one another. When $k=0$, it assesses exact matches. For this study, we set $k=1$ (AAR1), meaning raters’ scores are treated as matching as long as they differ by no greater than 1.
@@ 5.4.1 Scores
Figure 3 illustrates the overall scoring agreement between all pairs of assessors.
@@
More concretely, human-human QWK and AAR1 are almost always higher than the corresponding human-LLM agreement. Similarly, LLM-LLM agreement exceeds human-LLM agreement in virtually all cases, with a much larger margin, suggesting that LLMs may resemble each other in scoring more closely than humans resemble each other. Criterion-level agreement between human/LLM assessors shows similar patterns, as shown in Fig. 4.
@@ 5.4.3 Score-Comment Interaction
As expected, the last column in Table 2 shows strongly negative score-comment correlations across both human- and LLM-generated assessments. The fact that these negative correlations are generally much stronger when measured with the number of identified problems suggests that it is a more fine-grained metric than comment length and also indicates the usefulness of our framework proposed in Section 4.
@@ 6.1 Re-examining Our Assumption about Feedback Comment Quality
Our proposed feedback comment quality evaluation framework assumes that the quality of a feedback comment is related to how well it identifies relevant writing problems of an assessed essay. The framework extracts and characterizes problems of assessed essays identified in comments to evaluate the specificity and helpfulness of these comments.
To assess this assumption, we adopt an LLM-as-a-judge approach Zheng et al. (2023), prompting openai-o1-mini-2024-09-12 (o1-mini, OpenAI et al., 2024b) to directly assess the specificity and helpfulness of a feedback comment, given the corresponding essay and assessment question on a 10-point scale. We do not define specificity and helpfulness to avoid injecting biases and choose all comments, generated by humans and LLMs, from one subjective criterion (C6: coherence or flow of ideas) and one objective criterion (C9: academic vocabulary) to balance our examination. We then calculate the corrections between these two scores produced by o1-mini and the number of different types of problems identified by our framework.
The results in Table 3 shows that the characteristics extractable from applying the framework correlate very well with the o1-mini-assigned specificity and helpfulness scores. In particular, the number of problems that mention specific essay parts and offer corrections appears to be overall stronger signals of specificity and helpfulness than the mere number of problems, which shows negligible correlations for comments from IM 1 or IM 2. This shows the potential of our framework in providing a more fine-grained and interpretable measurement of specificity and helpfulness levels of comments.
@@ 6.2 Reliability of LLM-generated Assessments
The results in Table 4 show that LLMs are capable of generating highly stable scores, with an AAR1 score at least 0.81 and mostly above 0.9 across all conditions. Their generated comments are also decently similar with BERTScore typically no lower than 0.67. A small-scale manual check and a correlation analysis performed in Appendix D further verify the validity of BERTScore.
@@ 7 Conclusion
This study provides evidence that LLMs can generate reasonably good and generally reliable multi-dimensional analytic assessments. Our findings highlight the promising role of LLMs in assessing academic English writing, especially for graduate-level literature reviews, which is a highly technical genre. In short, LLMs show strong pedagogical potential, benefiting both L2 learners and instructors for self-regulated learning or teaching assistance. We propose and validate a novel feedback comment quality framework to facilitate our analysis.
Looking ahead, future studies could further characterize and compare the writing problems identified by human- versus LLM-generated comments, offering deeper qualitative insights. Additionally, it would be valuable to develop a metric grounded in our proposed framework that can directly compare the relative quality of two sets of comments.

### v2

@@ Abstract
The paper explores the performance of LLMs in the context of multi-dimensional analytic writing assessments, i.e. their ability to provide both scores and comments based on multiple assessment criteria. Using a corpus of literature reviews written by L2 graduate students and assessed by human experts against 9 analytic criteria, we prompt several popular LLMs to perform the same task under various conditions. To evaluate the quality of feedback comments, we apply a novel feedback comment quality evaluation framework. This framework is interpretable, cost-efficient, scalable, and reproducible, compared to existing methods that rely on manual judgments. We find that LLMs can generate reasonably good and generally reliable multi-dimensional analytic assessments. We release our corpus and code [^1] https://github.com/jaaack-wang/multi-dimensional-analytic-writing-assessments. for reproducibility.
@@ A Case Study of L2 Graduate-Level Academic English Writing
Zhengxiang Wangthanks: Zhengxiang Wang was a research assistant at the University of Saskatchewan for the research project that led to the creation of the corpus.   Veronika Makarova   Zhi Li   Jordan Kodner   Owen Rambow Department of Linguistics & Institute for Advanced Computational Science, Stony Brook University Department of Linguistics, University of Saskatchewan zhengxiang.wang@stonybrook.edu
@@ 1 Introduction
Figure 1: Multi-dimensional analytic assessments relevant to the corpus used in the study, where each assessment contains a score and a comment.
In recent years, large language models (LLMs) have emerged as promising tools for self-regulated writing assessments among L2 learners. A growing number of studies (Chiang and Lee, 2023; Mizumoto and Eguchi, 2023; Han et al., 2024; Yancey et al., 2023, i.a.) have indicated the general usefulness of LLMs for automated writing assessments. Given their increasing use for this task, the following question remains understudied: can LLMs provide reasonably good multi-dimensional analytic writing assessments? We use the phrase “reasonably good” intentionally, given the open-ended nature of the task, particularly generating essay-level feedback comments. [^2] Strictly speaking, there are two types of feedback: quantitative feedback (scores) and qualitative feedback (comments). We use “feedback comments” to refer to qualitative feedback.
@@
    We provide empirical evidence that LLMs can generate reasonably good and generally reliable multi-dimensional analytic writing assessments. This is the primary goal of this study; we do not argue in favor of a specific LLM, nor do we advocate replacing humans with LLMs for this task.
@@
    We propose and validate a novel LLM-based framework, ProEval, for evaluating the quality of feedback comments. ProEval is time- and cost-efficient, scalable, and reproducible, compared to manual judgments. It is also interpretable and fine-grained, compared to direct quality ratings.
@@ 2 Related Work
Unlike prior AWE systems, LLMs can be prompted in natural language to jointly score and comment on a given essay. A growing body of research has explored the use of LLMs for assessing L2 writing. For AES, LLMs have been examined for holistic scoring (Mizumoto and Eguchi, 2023; Yancey et al., 2023; Wang and Gayed, 2024), discourse coherence scoring (Naismith et al., 2023), and multi-dimensional analytic scoring (Yavuz et al., 2024; Banno et al., 2024). For feedback comment generation, LLMs have been studied for generating corrective comments (Mizumoto et al., 2024; Song et al., 2024), holistic comments (Behzad et al., 2024a, b), and multi-dimensional analytic comments (Guo and Wang, 2024; Behzad et al., 2024a; Han et al., 2024). Stahl et al. (2024) is the only study we know of which explores LLMs jointly performing scoring and feedback comment generation, but holistically. Moreover, the ASAP [^4] https://www.kaggle.com/competitions/asap-aes corpus they use contains short essays by native speakers from Grade 7 to Grade 10 and has no human reference comments.
@@ 3 Corpus
The corpus is a result of a large research project conducted at the University of Saskatchewan, a Canadian public research university, in 2021 with an aim to examine the developmental trajectory of literature review writing skills among L2 graduate students. The project involved three rounds of a 5-unit online tutorial series conducted over the course of 2021, with each round lasting 13 weeks (see Appendix A for details). Participation was voluntary, with 31 participants completing all five writing tasks across all rounds, and 20 further students completing at least one task before withdrawing.
@@
The corpus has been used in our previous studies (Li et al., 2023a, b; Makarova et al., 2024), although it has not been made public until now. These three studies only use a subset of the corpus, namely essays written individually or those based on topics 1, 3, and 5.
Among these studies, Li et al. (2023a, b) focus on individual writing development without examining feedback comments, placing their work within English for Academic Purposes rather than AWE. While Makarova et al. (2024) explore ChatGPT’s ability to assess L2 academic writing, they only compare model output to averaged human scores and aggregated comments, lacking criterion-level analysis. Their analysis is limited to surface features (e.g., word count, type-token ratio, comment length) and does not consider different prompting conditions. In contrast, this study offers a broader evaluation using the full corpus, distinct methodologies, and a more fine-grained analysis, with no substantial overlap with our prior work.
@@
The corpus authors comprise a diverse group of L2 learners, representing a wide range of first languages and enrolled in graduate programs across various disciplines at multiple Canadian universities. Their English proficiency ranged from upper-intermediate to advanced, with an average score equivalent to IELTS band score [^5] https://ielts.org/take-a-test/your-results/ielts-scoring-in-detail. 7 based on conversions from various standardized English language tests. Scores varied from IELTS 6.5 to 8.5, with a standard deviation of 0.55.
To support their writing, authors received a curated bibliography for each writing task, designed to facilitate literature review writing while reducing the burden of bibliographic searches. Before submitting their final drafts for expert assessments, they participated in peer review (for topics 1, 3, and 5) or group collaboration (for topics 2 and 4). These two measures were intended to enhance the overall quality of the submitted essays.
@@
Most essays in the corpus were assessed by three (94.3%) or two (5.0%) independent human experts. As illustrated in Fig. 1, the assessments consist of scores on a 10-point scale and comments based on 9 analytic assessment criteria. While scores were required, comments were optional for the assessors. Six assessors with professional experience in English language teaching assessed at different stages of the research project. Table 1 provides basic information about them.
The 9 assessment criteria include: (C1) material selection; (C2) material integration and citation; (C3) quality of key components; (C4) logic of structure; (C5) content and clarity of ideas; (C6) coherence (flow of ideas) ; (C7) cohesion (use of connectors); (C8) grammar and sentence structure; and (C9) academic vocabulary. Comparatively, criteria C2, C8, and C9 are more technical and objective, since there are clearer rules and conventions governing proper citation practices, grammatical correctness, and appropriate academic word usage. In contrast, other criteria require more interpretive judgment, making them relatively more subjective in nature. See Table 7 in Appendix A.3 for details about these criteria.
@@ F Graduate RA 2, 3 1-5 106
4 ProEval: A Novel Feedback Comment Quality Evaluation Framework
@@
4.1 The Framework
To address the issues of manual judgment, we propose ProEval, an automatic problem-focused evaluation framework that evaluates the quality of a feedback comment in terms of its ability to effectively identify relevant writing problems within the assessed essay. As illustrated in Fig. 2 (left), the framework utilizes LLMs to extract problems identified in feedback comments and to characterize their specificity and potential helpfulness. Rather than prompting an LLM to directly rate comment quality, which raises concerns about the reliability of LLM-based evaluators Doddapaneni et al. (2024), our approach prioritizes transparency and interpretability by grounding evaluation in concrete, observable features.
More concretely, ProEval consists of the following three steps, with the first two steps automated by GPT-4o-2024-11-20 OpenAI et al. (2024a) and the last step by GPT-4-Turbo-2024-04-09 OpenAI (2023). See Appendix B for additional details and explanations.
@@ 4.1 The Framework
The extracted problems are further characterized along three dimensions: whether an extracted problem (1) points to a specific part of the essay, (2) includes any form of suggestion (general or specific), and (3) provides a concrete correction that can be directly applied to fix an identified problem. These classifications offer a quantifiable way to assess the specificity and potential helpfulness of related comments.
@@
The results show that both human- and LLM-provided corrections are highly relevant, with answers to those three questions being “Yes” typically above 90% time (see Table 10 in Appendix B.3). We thus focus on the Problem Classification results in the next two sections.
4.2 Validations of the Framework
The basic idea of ProEval is to break down a complex and inherently subjective evaluation task into multi-level subtasks that are easy for humans to verify and well-suited for LLMs to perform. To validate that, the first author and a paid graduate student in Linguistics (native speaker) first annotated some held-out samples for training and developing the annotation guidelines. Each then independently annotated at least another 200 samples containing human- and LLM-generated comments or problems for Problem Extraction and Problem Classification. Afterward, they met to resolve disagreements before the inter-annotator agreement (IAA) was calculated.
We measure IAA using Cohen’s Kappa. As is known Feinstein and Cicchetti (1990), Cohen’s Kappa can provide misleading values with highly imbalanced class distributions. We therefore also provide exact match rates which have not been corrected for random agreement. Fig. 2 (right) shows that the IAA is typically high. When the Cohen’s Kappa is low due to class imbalance (i.e., problems being incorrectly or not extracted is uncommon or rare and nearly all extracted problems contain a suggestion), the exact match rates are high. LLM task performance, evaluated based on the resolved annotations, is also notably high (e.g., 0.92 F1 for Problem Extraction and at least 87% accuracy for the classification tasks in Problem Classification).
@@ 5.2 Baselines
Given the open-ended nature of the task, we compare raw assessments produced across individual assessors to understand the assessment patterns and behaviors of humans and LLMs. For a more robust statistical analysis, we only consider raw assessments made by assessors B, C, and F, since the essays they each assessed and co-assessed both cover at least half of the corpus (at least 78 essays between assessors C and F). See Table 11 in Appendix C.1 for exact numbers of essays all assessor pairs (including LLM assessors) co-assessed.
@@ 5.3 Evaluation of Scores
AAR measures the percentage of scores (from two raters) that lie within a specified threshold $k$ of one another. When $k=0$, it assesses exact matches. For this study, we set $k=1$ (AAR1), meaning raters’ scores are treated as matching or equivalent as long as they differ by no greater than 1.
@@ 5.4.1 Scores
Fig. 3 illustrates the overall scoring agreement between all pairs of assessors.
@@
More concretely, human-human QWK and AAR1 are almost always higher than the corresponding human-LLM agreement. Similarly, LLM-LLM agreement exceeds human-LLM agreement in virtually all cases, with a much larger margin, suggesting that LLMs may resemble each other in scoring more closely than humans resemble each other. This may be attributed to the substantial overlap in LLM training data, in contrast to the broader variability in human linguistic experiences, which contributes to greater divergence in human scoring patterns. Criterion-level agreement between human/LLM assessors shows similar patterns, as shown in Fig. 4.
@@ 5.4.3 Score-Comment Interaction
As expected, the last column in Table 2 shows strongly negative score-comment correlations across both human- and LLM-generated assessments. The fact that these negative correlations are generally much stronger when measured with the number of identified problems suggests that it is a more fine-grained metric than comment length and also indicates the usefulness of our framework (ProEval) proposed in Section 4. See Fig. 6 in Appendix C.4 for full results of the correlations.
@@ 6 Further Analyses
This section reexamines the assumption underlying our proposed feedback comment quality evaluation framework, i.e., ProEval, and evaluates the reliability of LLM-generated assessments.
@@ 6.1 Re-examining Our Assumption about Feedback Comment Quality
ProEval assumes that the quality of a feedback comment is related to how well it identifies relevant writing problems of an assessed essay. The framework extracts and characterizes problems of assessed essays identified in comments to evaluate the specificity and helpfulness of these comments.
To assess this assumption, we adopt an LLM-as-a-judge approach Zheng et al. (2023), prompting openai-o1-mini-2024-09-12 (o1-mini, OpenAI et al., 2024b) to directly assess the specificity and helpfulness of a feedback comment, given the corresponding essay and assessment question on a 10-point scale. We do not define specificity and helpfulness to avoid injecting biases and choose all comments, generated by humans and LLMs, from one subjective criterion (C6: coherence or flow of ideas) and one objective criterion (C9: academic vocabulary) to balance our examination. We then calculate the average Spearman rank correlations between these two scores produced by o1-mini and the number of different types of problems identified by ProEval under varying conditions.
The results in Table 3 shows that the characteristics extractable from applying ProEval correlate very well with the o1-mini-assigned specificity and helpfulness scores. In particular, the number of problems that mention specific essay parts and offer corrections appears to be overall stronger signals of specificity and helpfulness than the mere number of problems, which shows negligible correlations for comments from IM 1 or IM 2. This shows the potential of ProEval in providing a more fine-grained and interpretable measurement of specificity and helpfulness levels of comments.
@@ 6.2 Reliability of LLM-generated Assessments
The results in Table 4 show that LLMs are capable of generating highly stable scores, with an AAR1 score at least 0.81 and mostly above 0.9 across all conditions. Their generated comments are also decently similar with BERTScore typically no lower than 0.67. A small-scale manual check and a correlation analysis performed in Appendix D further verify the validity of BERTScore in measuring comment similarity.
@@ 7 Conclusion
This study provides evidence that LLMs can generate reasonably good and generally reliable multi-dimensional analytic assessments. Our findings highlight the promising role of LLMs in assessing academic English writing, especially for graduate-level literature reviews, which is a highly technical genre. In short, LLMs show strong pedagogical potential, benefiting both L2 learners and instructors for self-regulated learning or teaching assistance. We propose and validate a novel problem-focused evaluation framework, namely ProEval, to facilitate our analysis. Our stduy demonstrates that ProEval is time- and cost-efficient, scalable, and reproducible, compared to manual judgments. It is also interpretable and fine-grained, compared to direct quality ratings.
Looking ahead, future studies could further characterize and compare the writing problems identified by human- versus LLM-generated comments, offering deeper qualitative insights. Additionally, it would be valuable to develop a metric grounded in our proposed framework that can directly compare the relative quality of two sets of comments. We release our corpus to support continued research in this area.